A ball with mass 350 g is moving horizontally with a speed 12 m s-1


, hits a wall and

rebound at 15 m s-1 within 0.1 s. Calculate the impulse and the magnitude of force by

wall acting on the ball.


1
Expert's answer
2021-08-19T14:23:12-0400
"p=2mv=2(0.35)(12)=8.4\frac{kg\cdot m}{s}\\F=\frac{8.4}{0.1}=84\ N"